A 1500.0kg car driving at 15.0m/s[N] strikes a 3500.0kg truck travelling at 25.0m/s[E]. if the cars are tangled together during the collision determine the final velocity of both vehicles.


1
Expert's answer
2022-04-29T12:31:16-0400

"p_1=m_1v_1=22500,"

"p_2=m_2v_2=87500,"

"p=\sqrt{p_1^2+p_2^2}=90347~\frac{kg\cdot m}s."